The water drops are falling at regular intervals from the tap which is h m above the ground the third drop is about to leave the tap when first drop touches the ground  if the distance between the second drop and the first drop is 3.75 m then what is h
 

Grade:9

1 Answers

Arun
5 years ago
 

Height of tap = 5m and (g) = 10 m/sec2.

For the first drop, 5 = ut +  gt 2

= (0 × t) +  × 10 t2 v= 5tor t= 1 or t = 1.

It means that the third drop leaves after one second of the first drop. Or, each drop leaves after every 0.5 sec.

Distance covered by the second drop in 0.5 sec

 

        = (0.5)2 = 1.25m

        Therefore, distance of the second drop above the ground

= 5 - 1.25 = 3.75 m

 

You can check reversed